OpenStudy (anonymous):

Evaluate the given binomial coefficient (109/3)

OpenStudy (anonymous):

\[\binom{109}{3}\]?

OpenStudy (anonymous):

I have no idea how to do these problems.

OpenStudy (anonymous):

Yes, that is it.

OpenStudy (anonymous):

\[\frac{109\times 108\times 107}{3\times 2}\]

OpenStudy (anonymous):

else cheat http://www.wolframalpha.com/input/?i=109+choose+3

OpenStudy (anonymous):

How does one know which steps to follow to do this problem. Just to confirm, the answer would be the resulting number of that fraction?

OpenStudy (anonymous):

yes that is the result

OpenStudy (anonymous):

\[\binom{109}{3}=\frac{\overbrace{109\times 108\times 107}^{\text {three terms}}}{3!}\]

OpenStudy (anonymous):

So we know to go back three numbers (109, 108, 107) because it's over 3?

OpenStudy (anonymous):

Why do we multiply 3 by 2?

OpenStudy (anonymous):

the denominator is \(3!=3\times 2\)

OpenStudy (anonymous):

as another example \[\binom{10}{4}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}\]

OpenStudy (anonymous):

If the problem were, for instance (50/7), would it be (50x49x48x47x46x45x44/7x6x5x4x3x2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is a whole number in each case, so to compute, cancel first multiply last

OpenStudy (anonymous):

And (7/7) would be (7x6x5x4x3x2x1/7x6x5x4x3x2)? Basically I just want to confirm that the denominator just always counts down to 2?

OpenStudy (anonymous):

\(\binom{n}{n}=1\)

OpenStudy (anonymous):

yes, some people write a 1 there too, but that is silly because multiplying by one is like doing nothing

OpenStudy (anonymous):

like they might write \(4!=4\times 3\times 2\times1\)

OpenStudy (anonymous):

Gotcha. So the "denominator" will always count down to 2, and the "numerator" will count down the number of the denominator?

OpenStudy (anonymous):

doesn't count town to the number in the bottom counts down that many terms (like in the example you wrote)

OpenStudy (anonymous):

yeah i guess what you said is right if i interpret it correctly there is also a formula, but i wouldn't use it

OpenStudy (anonymous):

Yeah that's what I meant (the number of terms, not all the way down to that number). Thank you very much for explaining this all so clearly!

OpenStudy (anonymous):

yw